State if acid or base was present in the final mixed solution in Part B and show a sample calculation supporting your conclusion: 6 mL of 3.0 M HCL and 4 mL of 3.0 M NaOH
Answer #1

Neutralization reaction is

HCl + NaOH ............> NaCl + H2O

one mole NaOH is neutralized by one mole HCl.

we have

6 mL of 3.0 M HCL and 4 mL of 3.0 M NaOH.

thus

moles of HCl = volume of solution in L * molarity = 0.006 L * 3.0 mole / L = 0.018 mole HCl

and

moles of NaOH = volume of solution in L * molarity = 0.004 L * 3.0 mole / L = 0.012 mole

thus

excess HCl = (0.018 - 0.012) = 0.006 mole.

thus

acid was present in the final mixed solution.